# Understanding and Improving Information Preservation in Prompt Compression for LLMs

### v1

@@ E.1 Downstream Task Performance
Tables 12 and 13 present additional results for different variants of xRAG method on long-context datasets. Figure 4 presents downstream tasks performance of different variants of xRAG in terms of length of the context. The following xRAG variants are taken into account:
@@
Table 14: Grounding scores with respect to the source document/context for responses generated with different methods. All scores are an average across 5 random sets of 100 samples from the original evaluation set (standard deviation ($\sigma$) between sets is $<0.03$). “1st”: grounding score for the first claim detected in the response. “Avg”: average grounding score for all the claims in the response.
@@
Method xRAG Tokens HotpotQA HotpotQA* arXiv-sum. QuAC TriviaQA
@@
Table 16: Number of empty responses generated by xRAG models.
@@
Method xRAG Entire dataset 1-sent long samples 5-sents long samples 10-sents long samples
Tokens ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L
@@
Table 17: Results of information preservation by different variants of xRAG on the unseen (during training), out-of-domain examples. Performance is reported in terms of ROUGE metrics computed between the original and reconstructed text. The examples are encoded into one token directly or split into multiple tokens, one per sentence.
Method xRAG Entire dataset 1-sent long samples 5-sents long samples 10-sents long samples
Token ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L
@@
Table 18: Results of information preservation by different variants of xRAG on the seen (during training), in-domain examples. The ROUGE metrics are calculated in the same way as for the unseen set in Table 17.
@@ E.2 Grounding
Tables 14 and 15 report on grounding scores of the base compression methods and our xRAG modifications. In each table we present columns: 1st – grounding score of the first claim detected in the response, and Avg – the average grounding score for all the claims in the response. All the grounding scores for xRAG models are reported for the non-empty responses.
@@ E.3 Information Preservation
To better understand the information preservation in the xRAG compression method, we further extend our evaluation with ROUGE metrics (Tables 17 and 18). We use the same seen (during training, in-domain samples) and unseen (during training, out-of-domain samples) datasets as in Section 4.3. The low ROUGE scores (less than 0.5) indicate the models are not able to reconstruct the original content and instead, they are paraphrasing it, emphasizing the need for a metric that goes beyond lexical matching. Nevertheless, the ROUGE evaluations also show that our xRAG modification (xRAG w/ Two-Step PT + FT)) increases more than 2x the ROUGE scores reaching ROUGE-1 values close to 0.4 on the sentence level (both on seen and unseen data). This shows that the new representations are able to capture more precisely the original content compared to the tokens from the reproduced xRAG model. In addition to entity preservation results for “unseen” samples discussed in Section 5, we provide results for “seen” samples in Figure 6. Both datasets contain 450 examples, out of which 150 samples are one-sentence long, 150 contain 5 sentences and 150 contain 10 sentences. Unseen examples are sampled from three evaluation datasets: HoptpotQA (Yang et al., 2018), QuAC (Choi et al., 2018), and TriviaQA (Joshi et al., 2017) (with 150 examples from each of these datasets).
Qualitative Analysis
To understand the type of information preserved, we conducted a qualitative analysis of reconstructed examples, examining 10 samples of varying lengths generated using 20 different prompts. Our investigation focused on prompts from the xRAG paraphrase training phase. The analysis confirmed our hypothesis that the xRAG tokens primarily capture the general topic of the compressed content but fail to retain specific details. This behavior is expected, as xRAG relies on a dense retrieval model trained for document similarity. Depending on the topic, the LLM can reconstruct the main entity described in the text, aligning with its strong performance on the TriviaQA dataset. However, numerical entities such as dates and numbers are often lost. Examples of reconstructed contexts can be found in Table 23.
Data xRAG tokens Avg. PERSON GPE DATE CARD. ORG
Unseen 1 per sample 0.28 0.31 0.39 0.22 0.26 0.32
1 per sentence 0.19 0.21 0.25 0.14 0.13 0.24
Seen 1 per sample 0.25 0.16 0.38 0.24 0.38 0.21
1 per sentence 0.12 0.06 0.18 0.11 0.19 0.14
Table 19: Fraction of entities preserved in the reconstructed context using original xRAG model.

### v2

@@ E.1 Downstream Task Performance
Tables 14 and 15 present additional results for different variants of xRAG method on long-context datasets. Figure 4 presents downstream tasks performance of different variants of xRAG in terms of length of the context. The following xRAG variants are taken into account:
@@
Table 16: Grounding scores with respect to the source document/context for responses generated with different methods. All scores are an average across 5 random sets of 100 samples from the original evaluation set (standard deviation ($\sigma$) between sets is <0.03<0.03). “1st”: grounding score for the first claim detected in the response. “Avg”: average grounding score for all the claims in the response.
@@
Method Encodings HotpotQA HotpotQA* arXiv-sum. QuAC TriviaQA
@@
PISCO 8 per cont. 3 8 0 0 11
Table 18: Number of empty responses generated by soft prompting models.
@@
Method Encodings Entire dataset 1-sent long samples 5-sents long samples 10-sents long samples
ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L
@@
PISCO 8 per cont. 0.509 0.356 0.525 0.439 0.547 0.351 0.454 0.277
8 per sent. 0.555 0.409 0.509 0.415 0.608 0.437 0.549 0.374
Table 19: Results of information preservation by different variants of xRAG and PISCO on the unseen (during training), out-of-domain examples. Performance is reported in terms of ROUGE metrics computed between the original and reconstructed text. The examples are encoded into one token directly or split into multiple tokens, one per sentence.
Method Encodings Entire dataset 1-sent long samples 5-sents long samples 10-sents long samples
ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L ROUGE-1 ROUGE-L
@@
PISCO 8 per cont. 0.475 0.341 0.410 0.361 0.569 0.376 0.446 0.286
8 per sent. 0.553 0.421 0.420 0.371 0.633 0.460 0.606 0.431
Table 20: Results of information preservation by different variants of xRAG and PISCO on the seen (during xRAG training), in-domain examples. The ROUGE metrics are calculated in the same way as for the unseen set in Table 19.
@@ E.2 Grounding
Tables 16 and 17 report on grounding scores of the base compression methods and our xRAG modifications. In each table we present columns: 1st – grounding score of the first claim detected in the response, and Avg – the average grounding score for all the claims in the response. All the grounding scores for xRAG models are reported for the non-empty responses.
@@ E.3 Information Preservation
To better understand the information preservation in the xRAG compression method, we further extend our evaluation with ROUGE metrics (Tables 19 and 20). We use the same seen (during training, in-domain samples) and unseen (during training, out-of-domain samples) datasets as in Section 4.3. The low ROUGE scores (less than 0.5) indicate the models are not able to reconstruct the original content and instead, they are paraphrasing it, emphasizing the need for a metric that goes beyond lexical matching. Nevertheless, the ROUGE evaluations also show that our xRAG modification (xRAG w/ Two-Step PT + FT)) increases more than 2x the ROUGE scores reaching ROUGE-1 values close to 0.4 on the sentence level (both on seen and unseen data). This shows that the new representations are able to capture more precisely the original content compared to the tokens from the reproduced xRAG model. In addition to entity preservation results for “unseen” samples discussed in Section 5, we provide results for “seen” samples in Figure 5. Both datasets contain 450 examples, out of which 150 samples are one-sentence long, 150 contain 5 sentences and 150 contain 10 sentences. Unseen examples are sampled from three evaluation datasets: HoptpotQA (Yang et al., 2018), QuAC (Choi et al., 2018), and TriviaQA (Joshi et al., 2017) (with 150 examples from each of these datasets).